• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings/R. Zielinski

    Log entry time 15:42:45 on March 07, 2012

    Entry number 364291

    This entry is a followup to: 363098


    keyword=T2 Rate [late entry]

    Testing with cosmics, we found that there was a significant overlap of
    T1/T2 (~1% of events).
    Yesterday we swapped out the MLU but found that this did not solve the
    problem.  We suspected it might be the strobe, which was being formed
    with s2m coming first.  By switching the order of the strobe (so that s1
    came first), the T2 rate was as expected.  We suspect that the delay
    module used to delay s1 in the original configuration had some effect on
    the signal.  The strobe is still setup with s2m coming first, but the
    suspect delay module has been removed.  A cosmic test suggested that the
    problem is fixed, but we will continue to monitor with beam.

    A copy of this log entry has been emailed to: vasulk, rbziel, melissac